Suppose 1.5 cubic meter of oil is spilled into the ocean. Find the area of the resulting slick, assuming that it is one molecule thick, and that each molecule occupiesa cube 0.50 micro meters on a side.
Answer #1

.50 micrometers = 5x10-7 meters.

Volume of each molecule = (5x10-7 meters)3 = 1.25x10-19 cubic meters

1.5 cubic meters / 1.25x10-19 cubic meters = 1.2x1019 molecules

If it is 1 molecule thick, then each molecule has an area of:

(5x10-7 meters)2 = 2.5x10-13 meters2

2.5x10-13 meters2 x 1.2x1019 molecules = 3,000,000 meters2

answered by: Tionne